# CPathAgent: An Agent-based Foundation Model for Interpretable High-Resolution Pathology Image Analysis Mimicking Pathologists’ Diagnostic Logic

### v1

@@ Abstract
Recent advances in computational pathology have led to the emergence of numerous foundation models. However, these approaches fail to replicate the diagnostic process of pathologists, as they either simply rely on general-purpose encoders with multi-instance learning for classification or directly apply multimodal models to generate reports from images. A significant limitation is their inability to emulate the diagnostic logic employed by pathologists, who systematically examine slides at low magnification for overview before progressively zooming in on suspicious regions to formulate comprehensive diagnoses. To address this gap, we introduce CPathAgent, an innovative agent-based model that mimics pathologists’ reasoning processes by autonomously executing zoom-in/out and navigation operations across pathology images based on observed visual features. To achieve this, we develop a multi-stage training strategy unifying patch-level, region-level, and whole-slide capabilities within a single model, which is essential for mimicking pathologists, who require understanding and reasoning capabilities across all three scales. This approach generates substantially more detailed and interpretable diagnostic reports compared to existing methods, particularly for huge region understanding. Additionally, we construct an expert-validated PathMMU-HR², the first benchmark for huge region analysis, a critical intermediate scale between patches and whole slides, as diagnosticians typically examine several key regions rather than entire slides at once. Extensive experiments demonstrate that CPathAgent consistently outperforms existing approaches across three scales of benchmarks, validating the effectiveness of our agent-based diagnostic approach and highlighting a promising direction for the future development of computational pathology.
@@ 1 Introduction
Pathology serves as the gold standard for diagnosing numerous diseases, particularly cancer. Advances in digital pathology have enabled the digitization and computational analysis of histopathological slides, creating promising opportunities for AI-assisted diagnostics. Recently, various pathology foundation models have emerged to meet the growing demand for automated analysis in this field.
Current foundation models in computational pathology primarily fall into two categories (as shown in the left part of 1): encoder-based models like DINO[1; 2] or CLIP [3] that extract patch-level representations from Whole Slide Images (WSIs) followed by Multi-Instance Learning (MIL) [4; 5; 6] aggregation to perform diagnosis, and large multimodal models (LMMs) that process both visual data and text instructions to generate pathology reports or other outputs [7; 8; 9; 10]. Despite their impressive benchmark performance, these approaches share a critical limitation: they fail to incorporate the intermediate reasoning processes that expert pathologists employ, limiting their interpretability and clinical applicability.
In clinical practice, pathologists employ a methodical approach to slide examination, starting with low-magnification scanning (typically 2-4×) to assess tissue architecture and overall morphology. They then strategically navigate to regions of interest showing atypical features or diagnostic importance, progressively increasing magnification (10×, 20×, and 40×) to evaluate fine-grained histopathological features. Throughout this process, pathologists continuously shift between different magnifications and fields of view, integrating spatial and morphological information across multiple scales to formulate a comprehensive diagnosis. Current computational models bypass this sophisticated visual reasoning process, resulting in: (1) insufficient interpretability, as they cannot demonstrate visual navigation and multi-scale reasoning; (2) vulnerability to hallucinations and shortcut learning, potentially fixating on spurious correlations rather than relevant histological patterns; and (3) poor alignment with clinical workflows, undermining their credibility in diagnostic settings.
Drawing inspiration from recent breakthroughs in agent-based systems [11; 12; 13], which have demonstrated remarkable capabilities across diverse domains including strategic gameplay [14], web navigation [15], complex task automation [16], and multimodal content generation [17], we introduce CPathAgent, a novel framework that emulates the adaptive visual reasoning of expert pathologists. Unlike existing computational pathology models that operate as black-box classifiers or descriptors, CPathAgent implements a dynamic, perception-driven approach that mirrors pathologists’ hierarchical examination strategy in clinical practice. CPathAgent performs strategic operations: zooming in to examine cellular details, zooming out to maintain contextual awareness, and navigating across regions of interest, mimicking how pathologists scan slides at low magnification to identify suspicious areas before zooming in to examine fine-grained histological patterns before making diagnosis. This represents a fundamental paradigm shift in computational pathology from static, outcome-focused models to interactive, reasoning-based systems with greater clinical utility.
Specifically, CPathAgent mimics pathologists’ diagnostic workflows in a multi-stage process, as shown in the right part of Figure 1: (1) Global screening and region selection: CPathAgent initially analyzes a low-resolution WSI view to identify regions of interest (16000×16000 region at 40×) exhibiting suspicious histological features for detailed examination. (2) Navigation planning: For each prioritized region, CPathAgent plans a navigation path defining a sequence of view coordinates at varying magnifications, determining when to pan, zoom in, or zoom out based on the observed features within the huge region. (3) Multi-scale multi-view sequence reasoning: CPathAgent executes the planned navigation path by analyzing specific fields of view at different magnifications, then synthesizes observations across these views to perform step-by-step diagnostic reasoning.
Our work makes four significant contributions to advance agent-based computational pathology:
1) We introduce CPathAgent, which is a agent-based framework that emulates a pathologist’s diagnostic workflow through dynamic navigation planing and multi-magnification multi-view analysis, enabling interpretable, reasoning-based assessments.
@@
2) We develop PathMMU-HR², the first visual question answering benchmark specifically designed for high-resolution huge pathology image regions (16000×16000 pixels), rigorously validated by three professional pathologists to ensure clinical relevance and diagnostic accuracy.
@@
3) We introduce a multi-stage training strategy: An approach that equips CPathAgent with cognitive and reasoning capabilities across different scales, from patch-level details to whole-slide contexts, within a unified diagnostic framework.
@@
4) We conduct extensive evaluations across benchmarks at multiple scales, ranging from patches to large regions to WSIs, highlighting CPathAgent’s strong performance and potential clinical utility.
@@ 2 Related Work
2.1 Pathology Language-Vision Models
Recent years have witnessed substantial progress in developing foundation models specifically tailored for computational pathology. These models broadly fall into two categories: encoder-based models and large multimodal models (LMMs).
Encoder-based models primarily aim to extract patch-level representations from whole-slide images (WSIs). Early efforts in this area relied on ImageNet-pretrained convolutional neural networks (CNNs) [18; 19] as feature extractors, combined with multi-instance learning (MIL) techniques [4; 5; 6] to aggregate these features for slide-level prediction. More recently, the field has shifted toward pathology-specific foundation models that adopt self-supervised learning paradigms. Architectures such as DINO [1; 2] and CLIP [3], when trained on large-scale pathology datasets, have demonstrated superior feature extraction capabilities [20; 8; 21; 22; 23]. These improved representations, integrated via MIL or incorporated into pretrained WSI transformers, have significantly advanced the state of the art in WSI-level classification. Large multimodal models (LMMs) for pathology have emerged in parallel, enabling the processing of pathology images alongside text-based instructions to generate textual outputs. These models integrate visual and linguistic information through various designs. Patch-level MLLMs [7; 24; 8; 25; 26] adapt general vision-language frameworks like LLaVA [27] to interpret pathology image patches, while WSI-level models [9; 28; 29] aim to process entire whole-slide images for generating reports or answering diagnostic questions. Recent unified frameworks [10] attempt to handle both patches and WSIs within a single model architecture.
Despite these advances, current models typically process pathology inputs in a static and fixed manner, which contrasts with the dynamic, interactive diagnostic workflow of expert pathologists, who routinely navigate across regions of interest and examine tissues at multiple magnification levels, integrating contextual information over time to formulate a diagnosis.
@@ 2.2 Agent-based Models
Agent-based models represent an emerging paradigm in AI research, characterized by systems that can perceive their environment, make decisions, and take actions to achieve specific goals [30]. Recent advances in large language models (LLMs) have catalyzed significant progress in this domain, enabling more sophisticated reasoning and planning capabilities.Notable examples include WebGPT [15] and WebShop [31], which navigate web environments to complete information-seeking and shopping tasks; ReAct [32], which interleaves reasoning and action steps for improved task performance; AutoGPT [33] and AutoGen [34], which decompose complex tasks into manageable subtasks with minimal human supervision; and task-specific agents for code development [17], creative content generation [35], scientific research [36], and clinical decision-making and discovery that simulate physician workflows in hospital environments [37; 38; 39; 40]. These agent models typically incorporate several key components [11]: perception modules to process environmental inputs, reasoning modules to interpret observations and plan actions, and execution modules to implement decided actions. This architecture enables agents to adapt to dynamic environments and accomplish complex, multi-step tasks through iterative observation-reasoning-action cycles.
Although agent-based approaches have demonstrated success across various domains, their application to computational pathology remains largely unexplored. PathFinder [41] represents a notable early attempt at introducing agent-based systems for WSI analysis. The system segments WSIs into fixed 512×512 patches and employs a U-Net-based architecture to generate importance maps. To be specific, PathFinder feeds WSI thumbnails to generate initial importance maps, selects high-probability regions for description generation, and regenerate subsequent importance maps based on previous descriptions and thumbnails. While PathFinder introduces agent-based concepts to WSI analysis, it primarily operates through region selection at a fixed magnification level within predefined 512×512 patches. Pathologists, however, employ a more flexible diagnostic workflow with dynamic multi-scale navigation—beginning with low-magnification overviews and continuously adjusting zoom levels while moving across tissue regions based on morphological observations to integrate information across different scales.
To address this gap, our CPathAgent takes a step further by incorporating multi-scale navigation capabilities that mimic pathologist behavior. It autonomously decides where to examine, when to zoom in or out, and how to navigate across WSIs, closely mimicking the iterative diagnostic workflow of expert pathologists. Notably, CPathAgent provides real-time verbalization of its diagnostic reasoning, effectively simulating the internal thought processes that pathologists experience during slide examination (specific examples can be found in Appendix Figure 16 to 20). This approach enables enhanced interpretability through traceable reasoning paths and improved diagnostic performance.
@@ 2.3 Pathology Benchmarks and Datasets
Benchmark datasets play a crucial role in advancing computational pathology by providing standardized evaluation frameworks. Traditional whole-slide image (WSI) benchmarks like Camelyon16 [42] and TCGA [43] focused primarily on slide-level classification tasks. More recently, datasets like WSICaption[44] and WSI-VQA[45] have emerged, targeting whole-slide image description generation and question-answering tasks across entire slide images, respectively.
In parallel, several patch-level benchmarks have emerged for evaluating multimodal models in pathology. PathVQA [46] introduced the first visual question answering (VQA) tasks for pathology images, drawing on textbook image -caption pairs. QuiltVQA expanded this approach by extracting image-caption pairs from pathology lectures on YouTube, resulting in higher-quality QA pairs due to the educational nature of these materials. PathMMU [47] marks a more recent advance, offering a comprehensive benchmark across diverse pathology sources. Its key strength is the expert validation process, resulting in a professionally curated dataset that better reflects clinical practice.
Despite recent progress, a critical gap remains in existing benchmarks: they focus primarily on either small, localized patch-level tasks or broad whole-slide analysis, overlooking an essential intermediate scale, the huge region, which plays a central role in real-world diagnostic reasoning. In routine practice, pathologists typically begin by scanning the whole slide images to identify suspicious areas, and then concentrate their diagnostic attention on large, yet manageable regions containing the relevant pathology. It is often within these huge regions that key diagnostic decisions are made, as they provide sufficient context for interpretation.
To address this limitation, we introduce PathMMU-HR², a benchmark specifically designed for huge region analysis in pathology. By targeting this intermediate scale—situated between isolated patches and WSIs, PathMMU-HR² better aligns with the way pathologists actually examine and interpret tissue. It provides a more controlled, context-rich, and diagnostically meaningful setting for evaluating multimodal models, especially those incorporating agent-based reasoning across different scales. To ensure clinical relevance, we also engaged three board-certified pathologists to validate the dataset, making PathMMU-HR² an expert benchmark for intermediate-scale pathology evaluation.
@@ 3 Methods
In this section, we provide a comprehensive description of our proposed approach. We detail the construction of CPathAgent-Instruct, our instruction-tuning dataset used to train CPathAgent, the development of the PathMMU-HR2 benchmark dataset, and the specific training methodology implemented for the CPathAgent model.
3.1 Overview of CPathAgent Reasoning Workflow
@@
1) Global Screening: CPathAgent first scans the low-resolution thumbnail of a whole-slide image to identify and cluster suspicious regions. Uninformative areas are filtered out, while large diagnostically relevant regions (16000$\times$16000 pixels) are preserved for further analysis.
2) Navigation Planning: For each preserved huge region, the model plans a navigation path across spatial coordinates and magnification levels (e.g., 1$\times$, 3$\times$, 4.5$\times$). The goal is to plan sequentially examine areas that are most likely to yield diagnostic insights.
3) Multi-scale Multi-view Sequence Reasoning: Along the planned navigation path, the model observes each field of view, integrates visual features across scales and generates step-by-step reasoning descriptions similar to a pathologist’s thought process explaining the purpose of actions like zooming in and moving and finally synthesizes findings into a coherent diagnostic conclusion.

### v2

# envname-P envname#1

@@ Abstract
Recent advances in computational pathology have led to the emergence of numerous foundation models. These models typically rely on general-purpose encoders with multi-instance learning for whole slide image (WSI) classification or apply multimodal approaches to generate reports directly from images. However, these models cannot emulate the diagnostic approach of pathologists, who systematically examine slides at low magnification to obtain an overview before progressively zooming in on suspicious regions to formulate comprehensive diagnoses. Instead, existing models directly output final diagnoses without revealing the underlying reasoning process. To address this gap, we introduce CPathAgent, an innovative agent-based approach that mimics pathologists’ diagnostic workflow by autonomously navigating across WSI through zoom-in/out and move operations based on observed visual features, thereby generating substantially more transparent and interpretable diagnostic summaries. To achieve this, we develop a multi-stage training strategy that unifies patch-level, region-level, and WSI-level capabilities within a single model, which is essential for replicating how pathologists understand and reason across diverse image scales. Additionally, we construct PathMMU-HR², the first expert-validated benchmark for large region analysis. This represents a critical intermediate scale between patches and whole slides, reflecting a key clinical reality where pathologists typically examine several key large regions rather than entire slides at once. Extensive experiments demonstrate that CPathAgent consistently outperforms existing approaches across benchmarks at three different image scales, validating the effectiveness of our agent-based diagnostic approach and highlighting a promising direction for computational pathology.
@@ 1 Introduction
Pathology serves as the gold standard for diagnosing numerous diseases, particularly cancer. The advent of digital pathology has enabled the digitization and computational analysis of histopathological slides, opening transformative opportunities for AI-assisted diagnostics. Recently, numerous pathology foundation models have emerged to automate and enhance diagnostic workflows.
These foundation models (Figure 1, left) adopt two primary architectural paradigms. The first employs encoder-based models such as DINO[1; 2] or CLIP [3] to extract patch-level representations from WSIs, which are then aggregated via Multi-Instance Learning (MIL) [4; 5; 6] for diagnostic predictions. The second utilizes large multimodal models (LMMs) that jointly process visual data and text instructions to generate pathology reports or structured outputs [7; 8; 9; 10]. While these approaches achieve impressive benchmark performance, they share a fundamental mismatch with clinical practice. In real-world diagnostics, pathologists employ a systematic multi-scale examination strategy. They begin by scanning slides at low magnification (2-4×) to assess overall tissue architecture and identify regions of interest, then strategically navigating to atypical regions while progressively increasing magnification (10×, 20×, 40×) to examine fine-grained details. Throughout this process, pathologists continuously integrate observations across scales while conducting step-by-step diagnostic reasoning.
However, current models bypass this systematic sequential process by directly aggregating information across the WSI or processing visual inputs in a single forward pass, leading to two critical limitations. First, weak supervision: relying solely on slide-level labels without explicit guidance on diagnostically important regions or when to examine specific magnifications makes them vulnerable to shortcut learning [6], exploiting spurious correlations (staining artifacts, background patterns) rather than genuine pathological features. Second, lack of interpretability and verifiability: producing only final predictions without providing natural language descriptions of the intermediate reasoning prevents pathologists from understanding what the AI actually discovered and validating whether its findings are based on clinically relevant features.
To address these limitations, we introduce CPathAgent, a novel agent-based framework inspired by recent advances in AI agent systems that have achieved remarkable success from strategic gameplay [11] and web navigation [12] to task automation [13] and multimodal content generation [14; 15; 16; 17]. The key to their success lies in decomposing complex tasks into sequential reasoning steps with dynamic decision-making. Following this paradigm, CPathAgent bridges the gap between clinical practice and AI-assisted diagnosis by rethinking pathology diagnosis as adaptive visual reasoning. Unlike existing black-box models that treat diagnosis as single-shot classification, CPathAgent emulates pathologists’ examination strategy through WSI navigation with dynamic magnification adjustment and multi-view synthesis, conducting explicit step-by-step reasoning through natural language to enable interpretable and verifiable diagnostic assessments.
Specifically, CPathAgent performs a three-stage diagnostic workflow (Figure 1). In the global screening stage, CPathAgent analyzes a low-resolution WSI overview to identify suspicious regions that warrant detailed examination. In the navigation planning stage, CPathAgent plans an exploration strategy for each selected region, determining a sequence of view coordinates and magnification levels to investigate, mimicking how pathologists mentally map their examination before diving into details. Finally, in the multi-scale reasoning stage, the planned navigation path is executed through systematic analysis of fields of view at different magnifications (e.g., zooming in for cellular details, moving laterally across areas, zooming out for context), and synthesizes observations via step-by-step diagnostic reasoning to reach a conclusion. Overall, this work makes the following contributions:
1) CPathAgent framework: We propose an agent-based diagnostic system that performs diagnosis through strategic WSI navigation and multi-scale visual reasoning, mirroring pathologists’ systematic examination process while providing interpretable step-by-step rationales.
@@
2) PathMMU-HR² benchmark: We develop the first visual question answering benchmark specifically designed for high-resolution pathology image regions (16000×16000 pixels), rigorously validated by three professional pathologists to ensure clinical relevance and diagnostic accuracy.
@@
3) Multi-stage training strategy: We design a progressive training approach that equips CPathAgent with perceptual and reasoning capabilities across different image scales within a unified framework.
@@
4) Comprehensive evaluation: We conduct experiments at multiple scales, from patches and large regions to WSIs, highlighting CPathAgent’s strong performance and practical clinical utility.
@@ 2 Related Work
2.1 Pathology Foundation Models
Recent years have seen significant progress in developing foundation models for computational pathology, broadly categorized into encoder-based models and large multimodal models (LMMs).
Encoder-based models focus on extracting rich patch-level representations from WSIs. Early efforts in this area relied on ImageNet-pretrained CNNs [18; 19] as feature extractors, combined with multi-instance learning (MIL) [4; 5; 6] to aggregate these features for slide-level prediction. More recently, the field has shifted toward pathology-specific foundation models that adopt self-supervised learning paradigms. Architectures such as DINO [1; 2] and CLIP [3], when trained on large-scale pathology datasets, have demonstrated superior feature extraction capabilities [20; 8; 21; 22; 23]. These improved representations, integrated via MIL or incorporated into pretrained WSI transformers [20; 9], have significantly advanced the WSI-level classification performance. LMMs for pathology have emerged in parallel, enabling joint processing of pathology images and text instructions to generate textual diagnostic outputs. These models integrate visual and linguistic information through various designs. Patch-level LMMs [7; 24; 8; 25; 26] adapt general vision-language frameworks like LLaVA [27] for patch-level pathology image interpretation, while WSI-level models [9; 28; 29] aim to process entire WSIs for generating reports or answering diagnostic questions. Recent unified frameworks [10] attempt to handle both patches and WSIs within a single model architecture.
Despite these advances, current models typically process pathology inputs in a static manner, contrasting with expert pathologists who dynamically navigate regions of interest, examine tissues at multiple magnifications, and integrate contextual information to formulate diagnoses.
@@ 2.2 Agent-based Models
Agent-based models represent an emerging paradigm in AI research, characterized by systems that can perceive their environment, make decisions, and take actions to achieve specific goals [30]. Recent advances in large language models (LLMs) have catalyzed significant progress in this domain, enabling more sophisticated reasoning and planning capabilities. Notable examples include WebGPT [12] and WebShop [31], which navigate web environments to complete information-seeking and shopping tasks; ReAct [32], which interleaves reasoning and action steps for improved task performance; AutoGPT [33] and AutoGen [34], which decompose complex tasks into manageable subtasks with minimal human supervision; and task-specific agents for code development [14], creative content generation [35], scientific research [36], and clinical decision-making and discovery that simulate physician workflows in hospital environments [37; 38; 39; 40]. These agent models typically incorporate several key components [15]: perception modules to process environmental inputs, reasoning modules to interpret observations and plan actions, and execution modules to implement planned actions. This architecture enables agents to adapt to dynamic environments and accomplish complex, multi-step tasks through iterative observation-reasoning-action cycles.
Despite success in various domains, agent-based approaches remain largely unexplored in computational pathology. PathFinder [41] represents a successful early attempt at introducing agent-based systems for WSI analysis. The system segments WSIs into fixed 512×512 patches and employs a multi-agent framework that iteratively generates importance maps to select diagnostically relevant regions, produces natural language descriptions of these patches, and then regenerates subsequent importance maps conditioned on previous descriptions. However, PathFinder primarily operates through region selection at a fixed magnification level within predefined 512×512 patches, lacking the dynamic multi-scale navigation capability that characterizes pathologist workflows.
To address this gap, CPathAgent takes a step further by incorporating multi-scale navigation capabilities that mimic pathologist behavior. It autonomously decides where to examine, when to zoom in or out, and how to navigate across WSIs, closely mimicking the iterative diagnostic workflow of expert pathologists. Notably, CPathAgent provides verbalization of its diagnostic reasoning, effectively simulating the internal thought processes that pathologists experience during slide examination (specific examples can be found in Appendix Figures A12 to A16). This approach enables enhanced interpretability through traceable reasoning paths and improved diagnostic performance.
@@ 2.3 Pathology Benchmarks and Datasets
Benchmark datasets provide standardized evaluation frameworks crucial for advancing computational pathology. Traditional WSI benchmarks like Camelyon16 [42] and TCGA [43] focused primarily on slide-level classification tasks. More recently, datasets like WSICaption[44] and WSI-VQA[45] have expanded the scope to more complex tasks, including WSI description generation and VQA.
In parallel, several patch-level benchmarks have emerged for evaluating multimodal models in pathology. PathVQA [46] introduced the first VQA tasks for pathology images, drawing on textbook image-caption pairs. QuiltVQA [47] expanded this approach by extracting image-caption pairs from pathology lectures on YouTube, resulting in higher-quality VQA pairs due to the educational nature of these materials. PathMMU [48] marks a more recent advance, offering a comprehensive benchmark across diverse pathology sources validated by expert pathologists to ensure clinical relevance.
Despite this progress, existing benchmarks predominantly focus on either localized patch-level tasks or broad WSI analysis, overlooking a critical intermediate scale: the huge region. This scale is particularly important because it mirrors real-world diagnostic workflows. In routine practice, pathologists typically begin by scanning the WSIs to identify suspicious areas, then focus their diagnostic attention on these large, yet manageable regions. It is often within these huge regions that key diagnostic decisions are made, as they provide sufficient context for interpretation.
To address this limitation, we introduce PathMMU-HR², a benchmark specifically designed for huge region analysis in pathology. By targeting this intermediate scale between isolated patches and WSIs, PathMMU-HR² better aligns with how pathologists actually examine and interpret tissue. It provides a more controlled, context-rich, and diagnostically meaningful setting for evaluating multimodal models, especially those incorporating agent-based reasoning across different scales. To ensure clinical relevance, we also engaged three board-certified pathologists to validate the dataset, making PathMMU-HR² an expert benchmark for intermediate-scale pathology evaluation.
@@ 3 Methods
In this section, we detail the construction of CPathAgent-Instruct training data, the PathMMU-HR² benchmark, and the CPathAgent model architecture with its multi-stage training process.
3.1 Reasoning Workflow of CPathAgent
@@
1) Global Screening: To address the computational challenge of gigapixel WSIs, CPathAgent employs a coarse-to-fine region selection that mirrors pathologists’ systematic screening protocol.
[…]
2) Navigation Planning: For each preserved huge region $r_{i}$, the model generates a dynamic navigation plan as a sequence of viewing steps. Each step is specified by a tuple $(x,y,m,o)$, where $(x,y)$ denotes spatial coordinates in normalized space $[0,1]$, $m$ represents the magnification level relative to the original huge region (where $1.0\times$ corresponds to the full view), and $o$ describes the diagnostic focus for that position (i.e., what features need to be observed). The navigation plan $\mathcal{P}_{i}=\{(x_{1},y_{1},m_{1},o_{1}),\ldots,(x_{T},y_{T},m_{T},o_{T})\}$ is generated autoregressively, where each step is conditioned on the visual content of the region and WSI source information (e.g., lung, colon, etc.).
3) Multi-scale Multi-view Sequence Reasoning: Following the planned navigation path $\mathcal{P}_{i}$, the model receives the complete sequence of cropped images corresponding to all viewing steps $\{(x_{1},y_{1},m_{1},o_{1}),\ldots,(x_{T},y_{T},m_{T},o_{T})\}$ as input at once. Given this multi-view image sequence along with their associated diagnostic focus descriptions and spatial coordinates, the model performs holistic reasoning across the entire viewing trajectory in a first-person perspective similar to a pathologist’s examination process. The reasoning process systematically observes features across multiple scales and positions, establishes and refines diagnostic hypotheses by cross-referencing evidence, and maintains logical continuity throughout the navigation sequence. Finally, the model synthesizes all observations into a coherent pathological report that summarizes the diagnostic findings.
Illustrative examples, including global screening results, navigation path planning, and multi-scale reasoning outputs, are shown in Appendix Figure A12 to Figure A15. The prompts used for CPathAgent through these stages are provided in Appendix Figure A26 to Figure A28.